The ratio b/w the ages of two persons is 6:5.and sum of there ages is 77 then how many years later there ratio becomes 8:7?
Answer by lwsshak3(11628)

The ratio b/w the ages of two persons is 6:5.and sum of there ages is 77 then how many years later there ratio becomes 8:7
***
At present:
b+w=77
b=77-w
..
b/w=6/5
5b=6w
5(77-w)=6w
385-5w=6w
11w=385
w=385/11=35
b=6w/5=42
..
let x=years from now when b/w ratio becomes 8:7
(42+x)/(35+x)=8/7
7(42+x)=8(35+x)
294+7x=280+8x
x=294-280=14
ans:
The ratio 6:5 of the ages of two persons becomes 8:7 after 14 years